# Test-Time Training Provably Improves Transformers as In-context Learners

### v1

@@ 5 Analysis for General Covariance
(a) Losses after test-time-training as a function of $\gamma=k/d$ for optimal weights $\bm{W}^{*}$ based on two covariances and also $\bm{0}_{d\times d}$.
(b) Losses after test-time-training vs. $\gamma$ for optimal weights $\bm{W}^{*}$ with worst/best aligned covariances and $\bm{0}_{d\times d}$.
@@
Figure 2: Population losses in single and multi-step update scenarios. $\bm{\Sigma}_{\bm{x}}={\bm{I}}$, $\sigma^{2}=0$ and step sizes are optimal based on the formula in Theorem 5.3. $\bm{W}^{*}=\text{opt}$ and $\bm{W}^{*}=0$ represent the pre-trained matrix in (1) and zero (null) initialization, respectively. (a): Setting: $n=300$; $d=600$; $k$ changing between 00 and $4n$ with equal increments; $\bm{\Sigma}_{{\bm{\beta}},1}=\text{diag}({\bm{I}}_{250},0.5\cdot{\bm{I}}_{250});\bm{\Sigma}_{{\bm{\beta}},2}=\text{diag}(0.5\cdot{\bm{I}}_{250},{\bm{I}}_{250})$; ${\bm{\beta}_{\text{TT}}}=\begin{bmatrix}\bm{1}_{250}\;;0.5\cdot\bm{1}_{250}\end{bmatrix}$. Solid lines denote theoretical predictions which match the empirical (markers) results. (b): Setting: $n=250$; $d=500$; $k$ changing between 00 and $4n$ with equal increments; $\bm{\Sigma}_{{\bm{\beta}},1}=\text{diag}(1,\bm{0}_{499});\bm{\Sigma}_{{\bm{\beta}},2}=\text{diag}(0,{\bm{I}}_{499})$; ${\bm{\beta}_{\text{TT}}}=\begin{bmatrix}1\;;\bm{0}_{499}\end{bmatrix}$. (c): The figure illustrates the empirical results, which imply that with the initially optimal step size, a significant improvement can be gained by a single gradient step. Each line depicts a different reduce factor on the step size $\eta$ after each gradient step. Setting: $n=50$; $d=100$; $k=50\times d$; $\bm{\Sigma}_{\bm{\beta}}={\bm{I}}$; ${\bm{\beta}_{\text{TT}}}=\bm{1}_{100}$; $\bm{W}^{*}$ is optimal.
@@
[…]
@@
Let $n/d=\Theta(1)$, $\bm{\Sigma}_{\bm{x}}={\bm{I}}$, and $\sigma^{2}=0$. Suppose that the initial weight matrix $\bm{W}$ is a diagonal matrix whose eigenvalues are in the interval $[0,\frac{1}{n+1}]$. [^2] In the joint diagonalizability case, the restriction on $\bm{W}$ covers the space of $\bm{W}$ that is jointly diagonal with $\bm{\Sigma}_{\bm{\beta}}$ and $\bm{\Sigma}_{\bm{x}}$. Then, the optimal step size that minimizes the population loss given in (3) after the test-time training update is
@@
A more general and equivalent version of this result, which applies to any pair of jointly diagonalizable (not necessarily diagonal) covariance matrices $\bm{\Sigma}_{\bm{x}}$ and $\bm{\Sigma}_{\bm{\beta}}$ and to any $\bm{W}$ that is also jointly diagonalizable with them, can be found in Appendix C. The above approximations are robust and become exact as $n,d\to\infty$ while having a fixed ratio $n/d$; see Remark B.4. Theorem 5.3 characterizes the improvement due to test-time training for a rather general class of $\bm{W}$. It shows that, for fixed $\|\bm{W}\|_{F}^{2}$, the loss improvement increases monotonically with $A$. In other words, when the misalignment measure $A$ is larger, test-time training achieves greater performance gains. Consequently, considering the optimal pre-trained $\bm{W}^{*}$, if the eigenvalue spectrum of $\bm{\Sigma}_{\bm{\beta}}$ aligns with the entries of the task ${\bm{\beta}_{\text{TT}}}$ (i.e., larger eigenvalues match larger entries), then $A$ is larger and thus yields a bigger improvement.
@@
For the remainder of this section, we focus on the optimal pre-trained $\bm{W}^{*}$ constructed in Proposition 5.2, i.e. the minimum-Frobenius-norm solution of the pre-training loss (1). When we initialize the pre-trained model weight as $\bm{W}^{*}$, the loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ can be computed by combining Proposition 5.2 and Theorem 5.3. When we initialize the weights as $\bm{W}=\bm{0}_{d\times d}$, then the loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ is obtained by using the fact that ${\cal{L}}(\bm{0}_{d\times d})=\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$. By combining these results, we identify the phase transition point in the non-isotropic task covariance setting, which determines when test-time training outperforms training from scratch.
@@
We illustrate the findings of Corollary 5.4 in Figures 2(a) and 2(b) and make the following three observations: (i) The phase transition point is monotonic as a function of $c_{1}$, while keeping $c_{2}$ fixed, which implies that it is monotonic as a function of the misalignment term $A$. This means that, as we increase the alignment between the task covariance $\bm{\Sigma}_{\bm{\beta}}$ and the new task parameter ${\bm{\beta}_{\text{TT}}}$, the crossing point of the green and blue curves occurs later than that of the green and orange curves, as observed in Figure 2(a). (ii) The worst-aligned case over the set of all possible $(\bm{\Sigma}_{\bm{\beta}},{\bm{\beta}_{\text{TT}}})$ pairs is the one that satisfies ${\bm{\beta}_{\text{TT}}}^{\top}\bm{\Sigma}_{\bm{\beta}}{\bm{\beta}_{\text{TT}}}=0$. In this case, the corresponding $c_{1}$ is equal to $1$, which implies that the phase transition point in Corollary 5.4 is less than 00 as $c_{2}>0$. This means that training from scratch is always better than utilizing the test-time training for the worst-aligned case, which is depicted in Figure 2(b) as the orange curve. (iii) The best-aligned case over the set of all possible $(\bm{\Sigma}_{\bm{\beta}},{\bm{\beta}_{\text{TT}}})$ pairs is the one that satisfies ${\bm{\beta}_{\text{TT}}}^{\top}\bm{\Sigma}_{\bm{\beta}}{\bm{\beta}_{\text{TT}}}/\texttt{tr}(\bm{\Sigma}_{\bm{\beta}})=\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$. In this case, the corresponding $c_{1}$ scales on the order of $n^{-2}$ and $c_{2}$ on the order of $n^{-1}$, which in turn makes the phase transition point from Corollary 5.4 approximately $n$. When $k,n,$ and $d$ all grow while maintaining the ratios $\alpha=n/d$ and $\gamma=k/d$ fixed, this phase transition point approaches infinity. As a result, test-time training is always better than training from scratch in the best-aligned case, which is depicted in Figure 2(b) as the blue curve.
@@ 6 Experiments
(a) Accuracy of TabPFN model with and without test-time-training as a function of number of in-context samples $n$.
Figure 3: (a): Accuracy of TabPFN with and without test-time training. Setting: $k$ = 1000; $n$ changing between $1$ and $1000$. (b): Normalized loss after test-time-training with pre-training and zero initialization. Setting: $d=60$; $n=40$; $k$ changing between $64$ and $512$ in increments of $64$; $\sigma^{2}=0.01$; $\bm{\Sigma}_{\bm{\beta}}=\text{diag}(0.1\cdot{\bm{I}}_{25},0.5\cdot{\bm{I}}_{10},{\bm{I}}_{25})$; $\bm{\Sigma}_{\bm{x}}={\bm{I}}$. We sample ${\bm{\beta}_{\text{TT}}}$ from the distribution $\mathcal{N}(\bm{0},\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}})$ where $\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}}=\text{diag}({\bm{I}}_{25},0.5\cdot{\bm{I}}_{10},0.1\cdot{\bm{I}}_{25})$.
@@ 6.2 Experiments on TabPFN
We demonstrate that test-time training (TTT) can significantly reduce the context length required for a given performance on TabPFN v2 (Hollmann et al., 2025). Specifically, we evaluate the TabPFN v2 model on samples from a single dataset with different context lengths. For the TabPFN (blue curve in Figure 3(a)), we directly load the pre-trained model and vary the context window length during evaluation. In contrast, for TabPFN+TTT (orange curve in Figure 3(a)), we finetune the model using different context lengths with $k=1000$ samples. As the context length $n$ decreases, the samples are divided into $1000/n$ groups, where each group undergoes $50$ training iterations. The experiments are conducted on the T4 benchmark (Gardner et al., 2024), and we report the average performance across all datasets with enough training samples.
As shown in Figure 3(a), TabPFN with test-time-training allows the model with $200$ in-context samples to almost match the performance of the TabPFN without test-time-training with $1000$ in-context samples. This means that test-time training reduces the number of required samples for tabular classification by about $5$ times. A log-scaled version of Figure 3(a) is provided in Appendix D (Figure 4) for a clearer view of performance gains across different scales of $n$.
@@ 6.3 Experiments on GPT-2
We demonstrate how the pre-trained model compares against the scratch model under test-time training with distribution shift using GPT-2 architecture (Radford et al., 2019), as shown in Figure 3(b). Our results are in agreement with Figure 1(b).
Throughout the experiments, we consider the linear model with Gaussian data following Section 3. In order to obtain a pre-trained model, we sample multiple tasks from the distribution with task covariance $\bm{\Sigma}_{\bm{\beta}}$ and train the model from scratch on these tasks until convergence. In contrast, in the scratch setting, the model is initialized randomly without any pre-training. During test-time training, we evaluate the model on new task parameters sampled from the distribution $\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}}$, where the covariance structure is provided in the caption of Figure 3(b). We then apply test-time training to both models by varying the test-time training set size. The results are averaged over all the new tasks sampled from $\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}}$. Finally, we repeat the entire procedure three times using different random seeds and report the averaged results.
The results in Figure 3(b) show that pre-training is beneficial when the test-time training set size is small. However, as the training set size increases, training from scratch outperforms the pre-trained model, aligning with Figure 1(b).
@@ 7 Discussion
We have developed a theoretical framework to characterize how a single-step gradient update at test time enhances in-context learning. In the case of an isotropic covariance matrix, we analyzed the improvement in loss under test-time training as a function of the number of in-context samples ($n$), embedding dimension ($d$), and test-time training set size ($k$). We then extended our analysis to the non-isotropic covariance case, examining how the alignment between the new task parameter (${\bm{\beta}_{\text{TT}}}$) and the task covariance matrix ($\bm{\Sigma}_{\bm{\beta}}$) affects performance. Our results reveal a phase-transition threshold on $k$ in both isotropic and non-isotropic settings, beyond which zero initialization (training from scratch) can outperform pre-trained initialization. Empirical results align well with our theoretical findings, underscoring single-step test-time training as an efficient and effective approach.
As a future direction, one can investigate the analysis of TTT under a more general architecture and data model. Possible extensions to the architecture model might be to analyze the TTT with softmax attention or multilayer attention, whereas possible extensions to the data model might be to analyze K-Fold cross-validation instead of the current 1-Fold validation approach. Additionally, investigating TTT in unsupervised or semi-supervised in-context learning (ICL) settings presents another promising direction for extending this work.
@@ Appendix A Proofs for Section 3
^$\hat{{\cal{L}}}(\bm{W})\;=\;\sum_{i=n+1}^{n+k}\left(\bar{y}_{i}-\text{SM}({\bm{Z}_{i}},{\bm{W}})\right)^{2}=\sum_{i=n+1}^{n+k}(\bar{y}_{i}-\bm{\bar{x}}_{i}^{\top}\bm{W}{\bm{u}_{\text{context}}})^{2}=\|{\bm{y}_{\text{train}}}-{\bm{X}_{\text{train}}}\bm{W}{\bm{u}_{\text{context}}}\|_{2}^{2}.$
[…]

### v2

@@ 5 Analysis for General Covariance
(a) Losses after test-time-training as a function of $\gamma=k/d$ for optimal weights $\bm{W}^{*}$ based on two covariances and $\bm{W}=\bm{0}_{d\times d}$.
(b) Losses after test-time-training vs. $\gamma$ for optimal weights $\bm{W}^{*}$ with worst/best aligned covariances and $\bm{W}=\bm{0}_{d\times d}$.
@@
Figure 2: Population losses in single and multi-step update scenarios. $\bm{\Sigma}_{\bm{x}}={\bm{I}}$, $\sigma^{2}=0$ and step sizes are optimal based on the formula in Theorem 5.3. $\bm{W}^{*}=\text{opt}$ and $\bm{W}^{*}=0$ represent the pre-trained matrix in (1) and zero (null) initialization, respectively. (a): Setting: $n=300$; $d=600$; $k$ changing between 0 and $4n$ with equal increments; $\bm{\Sigma}_{{\bm{\beta}},1}=\text{diag}({\bm{I}}_{250},0.5\cdot{\bm{I}}_{250});\bm{\Sigma}_{{\bm{\beta}},2}=\text{diag}(0.5\cdot{\bm{I}}_{250},{\bm{I}}_{250})$; ${\bm{\beta}_{\text{TT}}}=\begin{bmatrix}\bm{1}_{250}\;;0.5\cdot\bm{1}_{250}\end{bmatrix}$. Solid lines denote theoretical predictions which match the empirical (markers) results. (b): Setting: $n=250$; $d=500$; $k$ changing between 0 and $4n$ with equal increments; $\bm{\Sigma}_{{\bm{\beta}},1}=\text{diag}(1,\bm{0}_{499});\bm{\Sigma}_{{\bm{\beta}},2}=\text{diag}(0,{\bm{I}}_{499})$; ${\bm{\beta}_{\text{TT}}}=\begin{bmatrix}1\;;\bm{0}_{499}\end{bmatrix}$. (c): The figure illustrates the empirical results, which imply that with the initially optimal step size, a significant improvement can be gained by a single gradient step. Each line depicts a different reduce factor on the step size $\eta$ after each gradient step. Setting: $n=50$; $d=100$; $k=50\times d$; $\bm{\Sigma}_{\bm{\beta}}={\bm{I}}$; ${\bm{\beta}_{\text{TT}}}=\bm{1}_{100}$; $\bm{W}^{*}$ is optimal.
@@
Note that the eigenvalues of $\bm{W}^{*}$ are smaller than $1/(n+1)$ as $\bm{\Sigma}_{\bm{\beta}}$ is a positive semi-definite matrix. This condition ensures the stability of the linear attention model described in (5) as the magnitude of the SM output scales with the context length $n$ when $\bm{W}$ is fixed. Assuming a similar criterion for the set of $\bm{W}$, we provide the optimal step size parameter and the corresponding test-time training loss.
@@
Let $n/d=\Theta(1)$, $\bm{\Sigma}_{\bm{x}}={\bm{I}}$, and $\sigma^{2}=0$. Suppose that the initial weight matrix $\bm{W}$ is a diagonal matrix whose eigenvalues are in the interval $0,\frac{1}{n+1}]$. [^2] In the joint diagonalizability case, the restriction on $\bm{W}$ covers the space of $\bm{W}$ that is jointly diagonal with $\bm{\Sigma}_{\bm{\beta}}$ and $\bm{\Sigma}_{\bm{x}}$. Then, the optimal step size that minimizes the population loss given in (3) after the test-time training update is
@@
A more general and equivalent version of this result, which applies to any pair of jointly diagonalizable (not necessarily diagonal) covariance matrices $\bm{\Sigma}_{\bm{x}}$ and $\bm{\Sigma}_{\bm{\beta}}$ and to any $\bm{W}$ that is also jointly diagonalizable with them, can be found in Appendix˜C. The above approximations are robust and become exact as $n,d\to\infty$ while having a fixed ratio $n/d$; see Remark 4.3. Theorem 5.3 characterizes the improvement due to test-time training for a rather general class of $\bm{W}$. It shows that, for fixed $\|\bm{W}\|_{F}^{2}$, the loss improvement increases monotonically with $A$. In other words, when the misalignment measure $A$ is larger, test-time training achieves greater performance gains. Consequently, considering the optimal pre-trained $\bm{W}^{*}$, if the eigenvalue spectrum of $\bm{\Sigma}_{\bm{\beta}}$ aligns with the entries of the task $\bm{\beta}_{\text{TT}}$} (i.e., larger eigenvalues match larger entries), then $A$ is larger and thus yields a bigger improvement.
@@
(a) Accuracy of TabPFN v2 model with and without test-time-training as a function of number of in-context samples $n$.
Figure 3: (a): Accuracy of TabPFN v2 with and without test-time training. Setting: $k$ = 1000; $n$ changing between 11 and $1000$. (b): Normalized loss after test-time-training with pre-training and zero initialization. Setting: $d=60$; $n=40$; $k$ changing between $64$ and $512$ in increments of $64$; $\sigma^{2}=0.01$; $\bm{\Sigma}_{\bm{\beta}}=\text{diag}(0.1\cdot{\bm{I}}_{25},0.5\cdot{\bm{I}}_{10},{\bm{I}}_{25})$; $\bm{\Sigma}_{\bm{x}}={\bm{I}}$. We sample $\bm{\beta}_{\text{TT}}$} from the distribution $\mathcal{N}(\bm{0},\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}})$ where $\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}}=\text{diag}({\bm{I}}_{25},0.5\cdot{\bm{I}}_{10},0.1\cdot{\bm{I}}_{25})$.
For the remainder of this section, we focus on the optimal pre-trained $\bm{W}^{*}$ constructed in Proposition˜5.2, i.e. the minimum-Frobenius-norm solution of the pre-training loss (1). When we initialize the pre-trained model weight as $\bm{W}^{*}$, the loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ can be computed by combining Proposition 5.2 and Theorem 5.3. When we initialize the weights as $\bm{W}=\bm{0}_{d\times d}$, then the loss ${\cal{L}}_{TT}({\bm{W}_{\text{TT}}})$ is obtained by using the fact that ${\cal{L}}(\bm{0}_{d\times d})=\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$. By combining these results, we identify the phase transition point in the non-isotropic task covariance setting, which determines when test-time training outperforms training from scratch.
@@
We illustrate the findings of Corollary 5.4 in Figures 2(a) and 2(b) and make the following three observations: (i) The phase transition point is monotonic as a function of $c_{1}$, while keeping $c_{2}$ fixed, which implies that it is monotonic as a function of the misalignment term $A$. This means that, as we increase the alignment between the task covariance $\bm{\Sigma}_{\bm{\beta}}$ and the new task parameter $\bm{\beta}_{\text{TT}}$}, the crossing point of the green and blue curves occurs later than that of the green and orange curves, as observed in Figure 2(a). (ii) The worst-aligned case over the set of all possible $(\bm{\Sigma}_{\bm{\beta}},{\bm{\beta}_{\text{TT}}})$ pairs is the one that satisfies ${\bm{\beta}_{\text{TT}}}^{\top}\bm{\Sigma}_{\bm{\beta}}{\bm{\beta}_{\text{TT}}}=0$. In this case, the corresponding $c_{1}$ is equal to $1$, which implies that the phase transition point in Corollary 5.4 is less than 0 as $c_{2}>0$. This means that training from scratch is always better than utilizing the test-time training for the worst-aligned case, which is depicted in Figure 2(b) as the orange curve. (iii) The best-aligned case over the set of all possible $(\bm{\Sigma}_{\bm{\beta}},{\bm{\beta}_{\text{TT}}})$ pairs is the one that satisfies ${\bm{\beta}_{\text{TT}}}^{\top}\bm{\Sigma}_{\bm{\beta}}{\bm{\beta}_{\text{TT}}}/\texttt{tr}(\bm{\Sigma}_{\bm{\beta}})=\|{\bm{\beta}_{\text{TT}}}\|_{2}^{2}$. In this case, the corresponding $c_{1}$ scales on the order of $n^{-2}$ and $c_{2}$ on the order of $n^{-1}$, which in turn makes the phase transition point from Corollary 5.4 approximately $n$. When $k,n,$ and $d$ all grow while maintaining the ratios $\alpha=n/d$ and $\gamma=k/d$ fixed, this phase transition point approaches infinity. As a result, test-time training is always better than training from scratch in the best-aligned case, which is depicted in Figure 2(b) as the blue curve.
@@ 6.2 Experiments on TabPFN
We demonstrate that test-time training (TTT) can significantly reduce the context length required for a given performance on TabPFN v2 Hollmann et al. (2025). Specifically, we evaluate the TabPFN v2 model on the The Tremendous TabLib Trawl (T4) dataset Gardner et al. (2024) for a more comprehensive evaluation, which is a large-scale high-quality collection of tabular benchmarks. Following the official TabPFN v2 implementation and our theoretical setup, we select the datasets containing at least 1,250 samples (with 1,000 for training, using an 80–20 split), limit the number of classes to 10 by choosing the most frequent ones, and restrict datasets to 100 randomly selected features. We also convert regression tasks into 10-class classification tasks based on quantiles to maintain consistency across training and evaluation. For each selected dataset from T4, we evaluate TabPFN v2 with context length $n$ varying from 1 to 1000. For the TabPFN (blue curve in Figure 3(a)), we directly load the pre-trained model and vary the context window length during evaluation. In contrast, for TabPFN+TTT (orange curve in Figure 3(a)), we finetune the model using different context lengths with $k=1000$ samples. As the context length $n$ decreases, the samples are divided into $1000/n$ groups, where each group undergoes $50$ training iterations. We then report the average performance across all datasets with enough training samples.
In the previous sections, we have theoretically shown that TTT reduces the required in-context sample size for queries from a new task by performing only a single adaptation step, thus enhancing the efficiency of inference. As empirical corroboration of this, Figure˜3(a) illustrates that TabPFN with test-time-training allows the model with $200$ in-context samples to almost match the performance of the TabPFN without test-time-training with $1000$ in-context samples. This means that test-time training reduces the number of required samples for tabular classification by about $5$ times. It is worth emphasizing that the sample complexity benefit is more evident near the zero-shot regime as TTT helps the model memorize new task dynamics by improving the accuracy from 0.2 to 0.45. A log-scaled version of Figure˜3(a) is provided in Appendix D (Figure 4) for a clearer view of performance gains across different scales of $n$.
@@ 6.3 Experiments on GPT-2
We demonstrate how the pre-trained model compares against the scratch model under test-time training with distribution shift using GPT-2 architecture Radford et al. (2019), as shown in Figure˜3(b). Our results are in agreement with Figure˜1(b).
Throughout the experiments, we consider the linear model with Gaussian data following Section 3. In order to obtain a pre-trained model, we sample multiple tasks from the distribution with task covariance $\bm{\Sigma}_{\bm{\beta}}$ and train the model from scratch on these tasks until convergence. In contrast, in the scratch setting, the model is initialized randomly using the GPT-2 architecture without any pre-training. During test-time training, we evaluate the model on new task parameters sampled from the distribution $\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}}$, where the covariance structure is provided in the caption of Figure 3(b). We then apply test-time training to both models by varying the test-time training set size and updating all parameters of the GPT-2 model. The results are averaged over all the new tasks sampled from $\bm{\Sigma}_{{\bm{\beta}_{\text{TT}}}}$. Finally, we repeat the entire procedure three times using different random seeds and report the averaged results.
The results in Figure˜3(b) show that using a pre-trained model with TTT is more advantageous when the test-time training set size is small. However, as the training set size increases, training from scratch outperforms the pre-trained model, aligning with Figure 1(b).
@@ 7 Discussion
We have developed a theoretical framework to characterize how a single-step gradient update at test time enhances in-context learning. In the case of an isotropic covariance matrix, we analyzed the improvement in loss under test-time training as a function of the number of in-context samples ($n$), embedding dimension ($d$), and test-time training set size ($k$). We then extended our analysis to the non-isotropic covariance case, examining how the alignment between the new task parameter ($\bm{\beta}_{\text{TT}}$}) and the task covariance matrix ($\bm{\Sigma}_{\bm{\beta}}$) affects performance. Our results reveal a phase-transition threshold on $k$ in both isotropic and non-isotropic settings, beyond which zero initialization (training from scratch) can outperform pre-trained initialization. Empirical results align well with our theoretical findings, and together they underscore that single-step test-time training offers significant performance improvements at low computational costs. Overall, our findings highlight test-time training as a lightweight, supervised enhancement to standard in-context learning.
As a future direction, one can investigate the analysis of TTT under a more general architecture and data model. Possible extensions to the architecture model might be to analyze the TTT with softmax attention or multilayer attention, whereas possible extensions to the data model might be to analyze K-Fold cross-validation instead of the current 1-Fold validation approach. Additionally, investigating TTT in unsupervised or semi-supervised in-context learning (ICL) settings presents another promising direction for extending this work. We hope that our results will serve as a foundation for future research on the interaction of TTT and ICL methods.
@@ Appendix A Proofs for Section 3
$\hat{{\cal{L}}}(\bm{W})\;=\;\sum_{i=n+1}^{n+k}\left(\bar{y}_{i}-\text{SM}({\bm{Z}_{i}},{\bm{W}})\right)^{2}=\sum_{i=n+1}^{n+k}(\bar{y}_{i}-\bm{\bar{x}}_{i}^{\top}\bm{W}{\bm{u}_{\text{context}}})^{2}=\|{\bm{y}_{\text{train}}}-{\bm{X}_{\text{train}}}\bm{W}{\bm{u}_{\text{context}}}\|_{2}^{2}.$
The gradient of $\mathcal$$\hat${L}} w.r.t. $\bm{W}$ is then obtained by differentiating the summation: